find the two digit number whose ten digit is four less than its unit digits. If the orignial number is two more than three ties the sum of the digit. The answer is one of the following four answers: 48 or 15 or 26 or 37
I am stumped.

Answer by mananth(16949)

let x be in the tens place
y in units place
..
10x+y is the number.
..
y-4 =x
y-x=4.............1
..
10x+y= 3(x+y)+2
10x+y=3x+3y+2
7x-2y=2...........2
..
multiply equation1 by 2 and add equation 2 to it
2y-2x=8
..
2y-2x+7x-2y=8+2
5x=10
x=2 the tens place.
plug the value of x in the equation 1
y-x =4
y-2=4
y=6
..
The number is 26